OpenStudy (anonymous):

A transformer has 175 turnings on the primary coil and 3000 turnings on the secondary coil. If the output voltage is 4000 V, what is the input voltage?

OpenStudy (cybershadow):

Ok e(p) = primary coil voltage(input voltage) e(s) = secondary coil voltage(output voltage)=4000 V Ns =Number of turns of secondary =3000 Np = Number of turns of primary=175 therefore From the principle of the transformer e(s)/e(p) = Ns/Np implies 4000/e(p) = 3000/175 e(p) = 4000*175/3000 e(p) = 233.3 V(Answer) So thats the required input voltage Keep me posted if its the correct answer, i m pretty sure it is :)